OpenStudy (anonymous):

if 3 divided by X equal 6 then what does X equal

OpenStudy (anonymous):

18

OpenStudy (anonymous):

actually it's 2 :P sorry

OpenStudy (anonymous):

3/x=6 is also equal to 3/6=x

OpenStudy (aravindg):

3/x=6 then x/3=1/6 now multiply by 3 on both sides to get x

OpenStudy (anonymous):

Thanks you guys helped me lots

OpenStudy (aravindg):

yw :)